# Tricritical Kibble–Zurek Scaling in Rydberg Atom Ladders

### v1

@@ Abstract
which provides a direct probe of the subleading exponent. It is known previously that the Binder ratio works the best in the ordered phase but scales not as well in the disordered phase [21]. To remedy this we introduce a disordered Binder ratio $\widetilde{U}=1-\frac{1}{3}\langle\widetilde{M}^{4}\rangle/\langle\widetilde{M}^{2}\rangle^{2}$. Here the disorder parameter is defined as
@@
where $\operatorname{SWAP}$ interchanges the atoms of each rung. The operator can be equivalently written as $\widetilde{M}=\operatorname{SWAP}_{1}(I_{2}+\operatorname{SWAP}_{2}(\cdots+\operatorname{SWAP}_{L-1}(I_{L}+\operatorname{SWAP}_{L})\cdots))$, which facilitates an efficient matrix-product operator (MPO) calculation. Unlike the ordered Binder ratio, which we choose the peak, the disordered Binder ratio exhibits a valley when tangentially passing through the tricritical point, as shown in End Matter Fig. 4(a). The disordered Binder valleys are collected, and the data collapse is shown in Fig. 2(d), which is fully consistent with the scaling relation in Eq. (4).
@@
A dynamical analog of Zamolodchikov’s $c$-theorem.— Quantum dynamics can be roughly classified by its characteristic time scale, with adiabaticity and quench the two opposite limits. In the case of KZ, a natural parameter is the ramping speed $s$, and $s$ is generally assumed to be small so that the physics is well controlled by the critical point. In previous discussions, this is also the regime we focused on. Now we go slightly beyond and focus on an intermediate-speed ramping.
@@
Heuristically, upon inspecting the phase diagram in Fig. 1(c) with a 0d tricritical point and a 1d critical line, we expect only the neighboring region of the tricritical point to be controlled by its universal property, and as one moves far away the Ising/Potts criticality takes over. The argument can be made more precise by examining the RG flow around the tricritical point, see Fig. 3(a). One sees that far away from the tricritical point, the RG flow generically takes a detour near the critical line, thus inheriting criticality from the latter [77, 78]. Returning to dynamics, it follows that only with a slow enough ramping can the tricritical exponents be observed. As the ramping speed increases, there should be a crossover to Ising/Potts exponents. This is fully supported by numerical simulations with rampings of slow to intermediate speed, as shown in Fig. 3(b, c). In the numerics we choose a sweet-spot direction of ramping, along which criticality extends the farthest in the ground state [66].
@@
These results are reminiscent of the celebrated Zamolodchikov’s $c$-theorem. The theorem dictates that along an RG flow the central charge $c$ has to decrease. This physics is again well represented by the case of tricritical points, where two different criticalities come into play, and is best illustrated by the RG flow in Fig. 3(a). In the dynamical setting, we find that the observed central charge $c$ decreases when we increase the ramping speed for reasons delineated above. We therefore argue that our findings amount to a dynamical analog of the $c$-theorem.
@@
Conclusion and discussion.— In this Letter, we systematically investigate the KZ scaling near tricritical points, using two- and three-leg Rydberg ladders. Through both conventional and tangential ramping protocols near the tricritical point, we accurately determine the critical exponents associated with the two relevant operators, in close agreement with their CFT predictions at the exact tricriticality. Additionally, we uncover a dynamical analog of Zamolodchikov’s $c$-theorem in an intermediate-speed ramping regime.
@@
We anticipate an imminent experimental realization of our proposals, particularly for the two-leg model which only requires a planar geometry and harbors TCI criticality. Looking forward, the TCI is closely related to spacetime supersymmetry—a long-sought concept in theoretical physics. In our setup, the critical exponents $\nu$ and $\nu^{\prime}$ correspond to the bosonic sector of the two lowest weight symmetric fields, $\epsilon$ and $\epsilon^{\prime}$. These fields exhibit power-law correlations, $\langle\epsilon_{i}\epsilon_{j}\rangle\sim|i-j|^{-2\Delta_{\epsilon}}$, with $\Delta_{\epsilon}=2-1/\nu$ (and similarly for $\epsilon^{\prime}$, giving $\Delta_{\epsilon^{\prime}}$). Notably, $\Delta_{\epsilon^{\prime}}-\Delta_{\epsilon}=1/\nu-1/\nu^{\prime}=1$. This relationship is no coincidence: the fermionic mode $\psi$ is linked to these fields through the supersymmetry generator, resulting in the relation $\Delta_{\psi}=\Delta_{\epsilon}+1/2=\Delta_{\epsilon^{\prime}}-1/2$, illustrating the structure of fermion bridging the two bosons. Hence, our protocol offers an effective method of corroborating this nontrivial ramification of emergent supersymmetry. On the other hand, directly measuring the corresponding fermion $\psi$ dynamically remains an open question, as it would likely require ramping fermionic operators [79, 47]. This introduces an interesting direction for future research, which we intend to explore further.
@@
Note added.— We thank the authors of [80] for sharing their preprint with us prior to its posting on arXiv. Similarly, they investigate the KZ mechanism at a supersymmetric tricritical point with two relevant directions. Although their model differs from ours, the same universality emerges, resulting in overlapping and mutually corroborating findings.
@@
Acknowledgments.— We are grateful to Hui Zhai for collaboration on relevant projects and stimulating ideas. We thank Wenjun Zhang for helpful discussions on the experimental details. H.W. is supported by China Postdoctoral Science Foundation under Grant No. 2024M751609 and Postdoctoral Fellowship Program of CPSF under Grant No. GZC20231364. C.L. is supported by Tsinghua University Dushi program. The DMRG and TEBD calculations are performed using the ITensor library [81].
Supplementary Materials for
@@
I Ground state phase diagrams
In this section, we discuss the ground state phase diagrams for the two- and three-leg ladders. In both cases, we use DMRG to obtain the ground state $|\psi\rangle$. We calculate the entanglement entropy $S(l)$ of a subsystem with length $l$,
$\rho(l)=\operatorname{Tr}_{L\smallsetminus l}|\psi\rangle\langle\psi|,\quad S(l)=\operatorname{Tr}\rho(l)\log\rho(l)$ (S1)
Then we fit the central charge of the state with $S(l)=\frac{c}{3}\sin(\pi l/L)+\mathrm{const}$. For each $\Omega$, a sweep over $\Delta$ finds the local maximum of $c$, which corresponds to the phase transition point. The maximum of $c$ along the phase boundary gives the tricritical position. Alternatively, the ground state Binder ratio crossing can be used to locate the phase boundary.
[…]
Figure S1: The phase diagram for two-leg (a) and three-leg (b) Hamiltonian, with tuning parameters $\Omega$ and $\Delta$.
II Expansion of lattice operators with conformal fields
For better understanding the universal properties of the tricritical point, it is helpful to relate lattice operators with conformal fields. From symmetry considerations, we have
$X=\langle X\rangle+a\epsilon+b\epsilon^{\prime}+\dots,\quad n=\langle n\rangle+c\epsilon+d\epsilon^{\prime}+\dots,$ (S2)
where $\epsilon,\epsilon^{\prime}$ are the leading and subleading symmetric operators with conformal dimensions $\Delta_{\epsilon},\Delta_{\epsilon^{\prime}}$. In this section, we explain how to extract these coefficients numerically.
The main analytical tool we use is the state–operator correspondence in CFT. Roughly speaking, this relates the energy eigenstates of a microscopic model to the primary fields of the field theory. Two important consequences follow. First, the energies of the lowest eigenstates are given by $E_{\phi}=\Delta_{\phi}+n$, where $\Delta_{\phi}$ is the conformal dimension of the field $\phi$ and $n$ labels the descendant level. Here we have shifted and rescaled the energy spectrum so that the ground state has zero energy and descendant states have unit energy intervals. We can thus identify the particular states $|\epsilon\rangle,|\epsilon^{\prime}\rangle$, see Fig. S2. Second, we have the matrix element $\langle\phi|\phi|0\rangle=(2\pi/L)^{\Delta_{\phi}}$. Combined with the expansion, Eq. (S2), we can extract the four coefficients of interest. The conformal dimensions and the numerical results are collected in Table S1. This is the data we used to plot the RG flow in Fig. 3(a) and Fig. 5(a).
Figure S2: The low-energy spectrum of TCI (a) and TCP (b). The states are identified by their conformal weights $(h,\bar{h})$, with conformal dimension (microscopically, energy) $\Delta_{\phi}=h+\bar{h}$ and conformal spin (microscopically, momentum) $s=h-\bar{h}$.
Tricritical $\Delta_{\epsilon}$ $\Delta_{\epsilon^{\prime}}$ $a$ $b$ $c$ $d$
TCI 1/5 6/5 $0.252$ $-0.299$ $0.347$ 0.073
 TCP 2/7 10/7 0.465 $-0.174$ 0.281 0.053
Table S1: Conformal dimensions and expansion coefficients of TCI and TCP.
Figure S3: (a) When we fix $\Omega=\Omega_{\mathrm{TCI}}$ and tuning $\Delta$, the data collapse is valid only for a narrow range. (b) If we choose the path with $\theta_{s}\approx 10^{\circ}$, the data collapse is valid for a much wider range.
@@ II Expansion of lattice operators with conformal fields
III Ground state scalings and the sweet-spot direction
@@
In this section, we discuss critical scalings near the tricritical point in the ground state. For a generic (i.e. not along the phase boundary), we expect $\xi_{\mathrm{GS}}\propto|\Delta-\Delta_{c}|^{-\nu}$. As one moves away from the tricritical point, the scaling generically does not hold. This can be seen from the finite-size scaling of the correlation length, expressed as $\xi_{\mathrm{GS}}/L=\widetilde{\mathcal{F}}((\Delta-\Delta_{\mathrm{TCI}})L^{1/\nu})$, where $\xi_{\mathrm{GS}}$ can be extracted from the disconnected two-point correlator. However, by numerically exploring various directions emanating from the tricritical point, we find a sweet-spot direction where the scaling behavior extends much farther than other directions, as shown in Fig. S3. We emphasize that this direction does not originate from universality, but offers a window that enables the study of criticality away from the critical point. Indeed, it is made good use of when we explore intermediate-speed ramping as detailed in the main text.

### v2

@@ Results
which provides a direct probe of the subleading exponent. It is known previously that the Binder ratio works the best in the ordered phase but scales not as well in the disordered phase [king2023quantum]. To remedy this, we introduce a disordered Binder ratio $\widetilde{U}=1-\frac{1}{3}\langle\widetilde{M}^{4}\rangle/\langle\widetilde{M}^{2}\rangle^{2}$. Here the disorder parameter is defined as
@@
where $\operatorname{SWAP}$ interchanges the atoms of each rung. The operator can be equivalently written as $\widetilde{M}=\operatorname{SWAP}_{1}(I_{2}+\operatorname{SWAP}_{2}(\cdots+\operatorname{SWAP}_{L-1}(I_{L}+\operatorname{SWAP}_{L})\cdots))$, which facilitates an efficient matrix-product operator (MPO) calculation. Unlike the ordered Binder ratio, which we choose the peak, the disordered Binder ratio exhibits a valley when tangentially passing through the tricritical point, as shown in Supplementary Note 3. The disordered Binder valleys are collected, and the data collapse is shown in Fig. 2d, which is fully consistent with the scaling relation in Eq. (4).
Dynamical Zamolodchikov’s $c$-theorem
@@
Quantum dynamics can be roughly classified by its characteristic time scale, with adiabaticity and quench the two opposite limits. In the case of KZ, a natural parameter is the ramping speed $s$, and $s$ is generally assumed to be small so that the physics is well controlled by the critical point. In previous discussions, this is also the regime we focused on. Now we go slightly beyond and focus on an intermediate-speed ramping.
@@
Heuristically, upon inspecting the phase diagram in Fig. 1c with a 0d tricritical point and a 1d critical line, we expect only the neighboring region of the tricritical point to be controlled by its universal property, and as one moves far away the Ising/Potts criticality takes over. The argument can be made more precise by examining the RG flow around the tricritical point, see Fig. 3a. One sees that far away from the tricritical point, the RG flow generically takes a detour near the critical line, thus inheriting criticality from the latter [wang2025kink, li2025random].
Returning to dynamics, it follows that the tricritical exponents can only be observed with sufficiently slow ramping. This is evident in Fig. 3b, where the upper-right corner of the figure shows a good data collapse when the horizontal axis is scaled by the tricritical exponent $\mu_{\text{TCI}}$. For faster ramping speeds, data for different system sizes spread out. However, if we rescale the horizontal axis by the critical exponent $\mu_{\text{Ising}}$, as shown in Fig. 3c, the data collapse is restored in the fast ramping regime. This indicates that in this regime, the emergent tricritical Ising is not observed, but the original Ising criticality manifests instead. This result holds generally for different ramping directions through the tricritical point. Specifically, by choosing a direction in which the tricriticality extends the furthest in the ground state (Supplementary Note 4), one might expect that only TCI would be observed in this regime. Nevertheless, we still observe Ising criticality dominating the dynamics. The appearance of two distinct critical exponents across different ramping speed ranges is a generic feature that reflects this rich renormalization group structure, as we also demonstrate for the three-leg case (see Supplementary Note 2).
@@
These results are reminiscent of the celebrated Zamolodchikov’s $c$-theorem. The theorem states that along an RG flow the central charge $c$ has to decrease. This physics is again well represented by the case of tricritical points in the ground state, where two different criticalities come into play, and is best illustrated by the RG flow in Fig. 3a. Although the KZ protocols do not directly measure the central charge, our observed crossover behavior still largely conforms to the same phenomenology, namely a flow from tricriticality (high $c$) to the Ising/Potts criticality (low $c$).
Discussion
@@
In this work, we systematically investigate the KZ scaling near tricritical points, using two- and three-leg Rydberg ladders. Through both conventional and tangential ramping protocols near the tricritical point, we accurately determine the critical exponents associated with the two relevant operators, in close agreement with their CFT predictions at the exact tricriticality. Additionally, we uncover a dynamical analog of Zamolodchikov’s $c$-theorem in an intermediate-speed ramping regime.
@@
We anticipate an imminent experimental realization of our proposals, particularly for the two-leg model which only requires a planar geometry and harbors TCI criticality. Looking forward, the TCI is closely related to spacetime supersymmetry—a long-sought concept in theoretical physics. In our setup, the critical exponents $\nu$ and $\nu^{\prime}$ correspond to the bosonic sector of the two lowest weight symmetry-preserving fields, $\epsilon$ and $\epsilon^{\prime}$. These fields exhibit power-law correlations, $\langle\epsilon_{i}\epsilon_{j}\rangle\sim|i-j|^{-2\Delta_{\epsilon}}$, with $\Delta_{\epsilon}=2-1/\nu$ (and similarly for $\epsilon^{\prime}$, giving $\Delta_{\epsilon^{\prime}}$). Notably, $\Delta_{\epsilon^{\prime}}-\Delta_{\epsilon}=1/\nu-1/\nu^{\prime}=1$. This relationship is no coincidence: the fermionic mode $\psi$ is linked to these fields through the supersymmetry generator, resulting in the relation $\Delta_{\psi}=\Delta_{\epsilon}+1/2=\Delta_{\epsilon^{\prime}}-1/2$, illustrating the structure of fermion bridging the two bosons. Hence, our protocol offers an effective method of corroborating this nontrivial ramification of emergent supersymmetry. On the other hand, directly measuring the corresponding fermion $\psi$ dynamically remains an open question, as it would likely require ramping fermionic operators [Bravyi2002, li2024supersymmetry]. This introduces an interesting direction for future research, which we intend to explore further.
@@
Note added: In a recent post [yin2025driven], the authors also investigate the KZ mechanism at a supersymmetric tricritical point with two relevant directions. Although their model differs from ours, the same universality emerges, leading to overlapping and mutually corroborating findings.
Methods
Ground state phase diagrams
We describe the method used to obtain the ground state phase diagrams for the two- and three-leg ladders. In both cases, we employ DMRG to compute the ground state $|\psi\rangle$.
For each Rabi frequency $\Omega$, we sweep the detuning $\Delta$ to calculate the Binder ratio of the ground state under periodic boundary conditions (PBC), $U_{\text{GS}}(L,\Omega,\Delta)=1-\frac{1}{3}\langle\psi|M^{4}|\psi\rangle/\langle\psi|M^{2}|\psi\rangle^{2}$. For different system sizes $L$, the Binder curve $U_{\text{GS}}$ versus $\Delta$ crosses at a single point, which is used to determine the transition point. By collecting the transition detuning values along with their corresponding Rabi frequencies, we obtain the phase boundary.
Next, we identify the tricritical point by calculating the central charge $c$ along the phase boundary. The maximum value of $c$ identifies the tricritical point [li2024supersymmetry]. Specifically, we calculate the entanglement entropy along the phase boundary. The entanglement entropy $S(l)$ of a subsystem with length $l$ is defined as $S(l)=-\operatorname{Tr}\rho(l)\log\rho(l)$, where $\rho(l)=\operatorname{Tr}_{L\setminus l}|\psi\rangle\langle\psi|$ is the reduced density matrix obtained from the ground state $|\psi\rangle$ calculated under PBC. The central charge of the state is then obtained by fitting the entanglement entropy to [Vidal2003, Calabrese-Cardy_2004]
$S(l)=\frac{c}{3}\log\left[\sin\left(\frac{\pi l}{L}\right)\right]+\text{const}.$ (6)
Figure 4: Central charge $c$ vs Rabi frequency $\Omega$ for two-leg (a) and three-leg (b) cases, using cesium 70S and sodium 69S. Insets show data collapse with horizontal axis $hL^{1/\nu^{\prime}}$, validating the scaling proposal of Eq. (3).
[…]
For the three-leg ladder case, we also performed the same calculation with the same parameters of cesium 70S and sodium 69S as used in the two-leg case, by simply adding one more atom to each rung, as shown in Fig. 1b. The central charge result is similar to its two-leg cousin, except that the numerical values of the central charge and critical exponents follow those of Potts and tricritical Potts, see Fig. 4b. However, it is quite heavy to perform dynamical evolution with the full setup. Hence, to demonstrate the universality feature of the physics, we also propose an effective model, as detailed in the following section. The same universality is found, with the phase diagram presented in Supplementary Note 1.
The effective model for the two-leg ladder
Here, we present an effective description of the two-leg ladder when the blockade condition is well satisfied. In each rung, two spin-1/2 atoms form a Hilbert space of four states. If both atoms cannot be simultaneously excited due to the blockade, then only three states remain accessible, and each rung can be effectively described as a spin-1 mode. To be more specific, we label the atoms on the upper rung as $i=1,3,5,\ldots,2L-1$ and the lower rung atoms as $j=2,4,6,\ldots,2L$. Each rung is labeled by $I=1,2,3,\ldots,L$, which also labels the corresponding spin-1. In this representation, the Hamiltonian (1) can be written in terms of $S^{x}$ and $S^{z}$ operators, with the following correspondence:
$\Omega(\sigma^{x}_{i}+\sigma^{x}_{j})$ $\longleftrightarrow\sqrt{2}\Omega S^{x}_{I}$ (7)
$\Delta(n_{i}+n_{j})$ $\longleftrightarrow\Delta S^{z}_{I}$
$V_{1}^{\mathrm{ab}}(n_{i}n_{i+1}+n_{j}n_{j+1})$ $\longleftrightarrow\frac{V_{1}^{\mathrm{ab}}-V_{2}^{\mathrm{ab}}}{2}S^{z}_{I}S^{z}_{I+1}$
$V_{2}^{\mathrm{ab}}(n_{i}n_{j+1}+n_{i+1}n_{j})$ $\longleftrightarrow\frac{V_{1}^{\mathrm{ab}}+V_{2}^{\mathrm{ab}}}{2}(S^{z}_{I})^{2}(S^{z}_{I+1})^{2}$
where the attractive interaction between dual-species atoms is given by $V_{1}^{\mathrm{ab}}=C_{\mathrm{ab}}/r_{1}^{6}$ (upper-upper/lower-lower interaction) and $V_{2}^{\mathrm{ab}}=C_{\mathrm{ab}}/(r_{1}^{2}+r_{2}^{2})^{3}$ (upper-lower/lower-upper interaction). Longer-range interactions between $S_{I}^{z}$ and $S_{I+2}^{z}$ can also be included, with repulsive interaction strengths $V_{1}^{\mathrm{aa}}=C_{\mathrm{aa}}/(2r_{1})^{6}$ and $V_{2}^{\mathrm{aa}}=C_{\mathrm{aa}}/((2r_{1})^{2}+r_{2}^{2})^{3}$, and similar expressions for $V_{1}^{\mathrm{bb}}$ and $V_{2}^{\mathrm{bb}}$.
The same universality is observed in the ground state phase diagram, with the non-universal feature that the position of the tricritical point is shifted by approximately $5\%$ when only attractive effects are included. When longer-range repulsive interactions are also included, the position of the tricritical point shifts by approximately $0.6\%$ compared to the full range spin-$1/2$ model.
The effective model for the three-leg ladder
The effective Hamiltonian for the three-leg ladder reads
$H=\sum_{i}\frac{\Omega}{2}X_{i}-\Delta n_{i}-Jn_{i}n_{i+1}.$ (8)
In this effective model, by enforcing Rydberg blockade, each site has four states $|0\rangle$, $|a\rangle$, $|b\rangle$ and $|c\rangle$, corresponding to all atoms in the ground state or one of the atoms on a rung excited to the Rydberg state, respectively. $X=\sum_{\alpha=a,b,c}|\alpha\rangle\langle 0|+\mathrm{h.c.}$ couples the ground state and the Rydberg state as usual, and $n=\sum_{\alpha=a,b,c}|\alpha\rangle\langle\alpha|$ projects to the Rydberg states. Inter-rung attractive interactions are encompassed in a coupling constant $J$, which we set to 1 as the energy scale.
To properly work with the three-leg ladder, we need to redefine several physical quantities. The order parameter is defined by
$M_{3}=\sum_{i}n_{ia}+e^{2\pi\mathrm{i}/3}n_{ib}+e^{4\pi\mathrm{i}/3}n_{ic},$ (9)
and the disorder operator reads
$\widetilde{M}_{3}=\sum_{i}\prod_{j\geq i}\operatorname{CYCL}_{j},$ (10)
where $\operatorname{CYCL}=|0\rangle\langle 0|+|b\rangle\langle a|+|c\rangle\langle b|+|a\rangle\langle c|$ permutes the atoms along the rungs. A numerics-friendly version is $\widetilde{M}_{3}=\operatorname{CYCL}_{1}(I_{2}+\operatorname{CYCL}_{2}(\cdots+\operatorname{CYCL}_{L-1}(I_{L}+\operatorname{CYCL}_{L})\cdots))$. The Binder ratio can then be defined in the same way as described in the main text.
Numerical simulation of KZ dynamics
[…]
For the three-leg case, we use Hamiltonian (8) for TEBD simulations, and results are summarized in Supplementary Note 2.
Critical exponents and scaling dimensions
In the main text, we identified two critical exponents $\nu$ and $\nu^{\prime}$ near the tricriticality. In the CFT language, these exponents correspond to two symmetry-preserving relevant conformal fields, $\epsilon$ and $\epsilon^{\prime}$, in tricritical Ising/Potts CFT, with scaling dimensions $\Delta_{\epsilon}$ and $\Delta_{\epsilon^{\prime}}$, respectively.
The $\epsilon$ field is very similar to its counterpart at the parent critical line. It drives the system between disordered and ordered phases. For a shifted parameter $g$, the field couples to it with an additional action term compared to the critical point value: $S=\int\mathrm{d}x\,\mathrm{d}\tau\,g\cdot\epsilon$. Near criticality, scaling invariance is still preserved and action is unchanged under the following scaling transformations with ratio $b$:
$x\to b^{-1}x,\,\,\tau\to b^{-z}\tau,\,\,g\to b^{1/\nu}g,\,\,\epsilon\to b^{\Delta_{\epsilon}}\epsilon$ (11)
This immediately gives the relation between the critical exponents that can be measured in experiments and the scaling dimension of the coupled conformal field: $\Delta_{\epsilon}+1/\nu=1+z$ [cf. Table 1 and Table 2].
Tricritical $\Delta_{\epsilon}$ $\Delta_{\epsilon^{\prime}}$ $a_{1}$ $a_{2}$ $b_{1}$ $b_{2}$
TCI $1/5$ $6/5$ 0.252 0.299-0.299 0.347 0.073
TCP $2/7$ $10/7$ 0.465 0.174-0.174 0.281 0.053
Table 2: Conformal dimensions and expansion coefficients of TCI and TCP.
Figure 5: The low-energy spectrum of TCI (a) and TCP (b). The states are identified by their conformal weights $(h,\bar{h})$, with conformal dimension (microscopically, energy) $\Delta_{\phi}=h+\bar{h}$ and conformal spin (microscopically, momentum) $s=h-\bar{h}$.
The $\epsilon$ field, as the most relevant operator, is responsible for the spontaneous symmetry breaking. However, in the phase diagram spanned by two parameters, there is a direction that does not lead to either phase—namely, the direction parallel to the local phase boundary. In this direction, the perturbation involving $\epsilon$ is completely canceled, and the subleading field $\epsilon^{\prime}$ takes charge. We denote the angle of the phase boundary relative to the detuning direction as $\theta$, and the deviation in this direction from the tricritical point as $h_{\parallel}$. This angle $\theta$ can be determined by expanding lattice operators in terms of conformal fields, as illustrated in the next subsection. $h_{\parallel}$ refers to the strength of deviation from the tricritical point, with the equilibrium critical exponent defined as $\nu^{\prime}$. Since $h_{\parallel}$ couples to $\epsilon^{\prime}$, which contributes most significantly to the action, following the same argument gives $\Delta_{\epsilon^{\prime}}+1/\nu^{\prime}=1+z$. In the main text, we define $h=\Omega-\Omega_{\mathrm{TCI}}$ to control the deviation from the tricritical point, which has the relation $h=h_{\parallel}\sin(\theta)$. We use the dimension $h]=1/\nu^{\prime}$ in this sense, as verified by the central charge scaling along the phase boundary shown in the insets of Fig. 4.
With these equilibrium features near the tricritical point, we therefore propose a “tangential ramping” protocol that tangentially approaches the tricritical point. In this direction, before reaching the near-tricritical regime, the energy gap remains finite. Near the tricritical point, the KZ mechanism becomes relevant due to the gapless nature, with the scaling entirely controlled by $\epsilon^{\prime}$ and thus enabling measurement of $\mu^{\prime}$.
Expansion of lattice operators by conformal fields
In this section, we explain how we obtain the renormalization group flow at the tricritical point shown in Fig. 3a. We relate lattice operators $\sigma^{x}$, $n$ [cf. Eq. (1)] to conformal fields. From symmetry considerations, we have
$\sigma^{x}$ $=\langle\sigma^{x}\rangle+a_{1}\epsilon+a_{2}\epsilon^{\prime}+\dots$ (12)
$n$ $=\,\,\langle n\rangle\,+b_{1}\epsilon\,+b_{2}\epsilon^{\prime}+\dots$
We note that $\sigma^{x}$ and $n$ are coupled to $\Omega$ and $\Delta$, respectively. Once we determine these coefficients, and hence the relative strength between $\epsilon$ and $\epsilon^{\prime}$, we can establish (i) the direction of the phase boundary (where the $\epsilon$ component vanishes) and (ii) the generic RG flow trajectory in the phase diagram with parameters $\Omega$ and $\Delta$. We explain how to extract these coefficients numerically below.
The main analytical tool we use is the state–operator correspondence in CFT. Roughly speaking, this relates the energy eigenstates of a microscopic model to the primary fields of the field theory. Two important consequences follow. First, the energies of the lowest eigenstates are given by $E_{\phi}=\Delta_{\phi}+n$, where $\Delta_{\phi}$ is the conformal dimension of the field $\phi$ and $n$ labels the descendant level. Here we have shifted and rescaled the energy spectrum so that the ground state has zero energy and descendant states have unit energy intervals. We can thus identify the particular states $|\epsilon\rangle,|\epsilon^{\prime}\rangle$, see Fig. 5. Second, we have the matrix element $\langle\phi|\phi|0\rangle=(2\pi/L)^{\Delta_{\phi}}$. Combined with the expansion, Eq. (12), we can extract the four coefficients of interest. The conformal dimensions and the numerical results are collected in Table 2. This is the data we used to plot the RG flow in Fig. 3a. We also mention that the densest direction of the RG trajectory lines coincides with the DMRG phase boundary calculation (blue line, obtained using the Binder ratio), showing good agreement in determining the tangential direction.
Acknowledgements
@@
We are grateful to Hui Zhai for collaboration on related projects and stimulating ideas. We thank Wenjun Zhang for helpful discussions on the experimental details. We thank Shuai Yin for reading our manuscript and providing valuable feedback. This work is supported by National Natural Science Foundation of China under Grant No. 12504307 (C.L.). C.L. is also supported by Tsinghua University Dushi program. H.W. is supported by China Postdoctoral Science Foundation under Grant No. 2024M751609 and Postdoctoral Fellowship Program of CPSF under Grant No. GZC20231364. The DMRG and TEBD calculations are performed using the ITensor library [ITensor].
Author contributions
H.W., X.L., and C.L. contributed to developing the theory. H.W. performed the numerical simulations. All authors analyzed the data and wrote the manuscript.
Data availability
The data generated in this study have been deposited in the Zenodo database under identifier https://doi.org/10.5281/zenodo.17385416.
Code availability
The numerical codes used in this study are available from the corresponding author upon request.
Supplementary Information for
@@
Hanteng Wang, Xingyu Li, and Chengshu Li *
Institute for Advanced Study, Tsinghua University, Beijing, 100084, China
Supplementary Note 1: Ground state properties
We present the numerical results for the ground state phase diagram of both the two-leg and three-leg models in Supplementary Fig. S1. The phase boundary is determined by Binder ratio crossings, and the tricritical point is identified by the peak of the central charge, as shown in Fig. 4 in the main text.
Supplementary Note 2: More on numerical results for KZ dynamics
We present the numerical results of all three KZ protocols for the three-leg ladder in Supplementary Fig. S2 and Supplementary Fig. S3. All results show behavior nearly identical to those observed for the two-leg ladder, indicating that our proposals and findings are equally valid beyond the TCI case.
Supplementary Note 3: More on tangential ramping
Using phase boundary data, we can determine the tangential line for the tricritical point, denoted as $\theta_{t}$ for the angle. For the TCI, $\theta_{t}\approx 70^{\circ}$, and for the TCP, $\theta_{t}\approx 50^{\circ}$. These results are consistent with CFT calculations at the tricritical point, as shown in Fig. 3a and Supplementary Fig. S3a. With the tangential line determined, we prepare an initial state with zero Rabi frequency and then ramp both the Rabi frequency and detuning simultaneously. The coordinates of the ramping line on the phase diagram $(\Delta,\Omega)$ follow $\Delta=\Delta_{\mathrm{TC}}+st\cos\theta_{t}$, $\Omega=\Omega_{\mathrm{TC}}+st\sin\theta_{t}$, where $s$ is the ramping speed.
Using the disorder parameter, we plot how its corresponding Binder ratio $\widetilde{U}$ evolves over time as shown in Supplementary Fig. S2a. It starts at $2/3$ and then decreases. After tangentially touching the tricritical point, the Binder ratio exhibits a valley structure, which we use as the reference point for the subsequent scaling procedure. Figure 2d and Supplementary Fig. S2d show decent data collapse for this ramping protocol, allowing us to extract the subleading critical exponents $\mu^{\prime}$.
@@
Supplementary Note 4: Ground state scalings and the sweet-spot direction
@@
In this section, we discuss critical scalings near the tricritical point in the ground state. For a generic (i.e. not along the phase boundary), we expect $\xi_{\mathrm{GS}}\propto|\Delta-\Delta_{c}|^{-\nu}$. As one moves away from the tricritical point, the scaling generically does not hold. This can be seen from the finite-size scaling of the correlation length, expressed as $\xi_{\mathrm{GS}}/L=\widetilde{\mathcal{F}}((\Delta-\Delta_{\mathrm{TCI}})L^{1/\nu})$, where $\xi_{\mathrm{GS}}$ can be extracted from the disconnected two-point correlator. However, by numerically exploring various directions emanating from the tricritical point, we find a sweet spot direction where the scaling behavior extends much farther than other directions, as shown in Supplementary Fig. S4. We emphasize that this direction does not originate from universality, but offers a window that enables the study of criticality away from the critical point. Indeed, it is made good use of when we explore intermediate-speed ramping as detailed in the main text. Interestingly, along this direction where the ground state is controlled by a single scaling law, KZ dynamics reveals a crossover effect: at slow speeds, tricritical scaling is observed, while at faster speeds, Ising/Potts scaling is recovered. This differs from cases where dynamical scaling is inherited from equilibrium scaling. Our observed crossover is a purely dynamical phenomenon arising from the interplay between two relevant operators.
Supplementary Fig. S1: The phase diagram for two-leg (a) and three-leg (b) Hamiltonian, with tuning parameters $\Omega$ and $\Delta$.
Supplementary Fig. S2: a, Binder ratio $\widetilde{U}$ vs $(\Delta-\Delta_{\mathrm{TCP}})/\cos\theta_{t}\propto t$ defined in tangential case. The first valley is marked with a cross. b, KZ ramping across the TCP point. c, KZ ramping near the TCP point. d, Tangential KZ for TCP.
Supplementary Fig. S3: a, RG flow near the TCP. b, Slow ramping data collapse with $\mu_{\mathrm{TCP}}$. c, Fast ramping data collapse with $\mu_{\mathrm{Potts}}$.
Supplementary Fig. S4: a, When we fix $\Omega=\Omega_{\mathrm{TCI}}$ and tuning $\Delta$, the data collapse is valid only for a narrow range. b, If we choose the path with $\theta_{s}\approx 10^{\circ}$, the data collapse is valid for a much wider range.